# Did I Vet You Before? Assessing the Chrome Web Store Vetting Process through Browser Extension Similarity

### v1

@@ 6.2. No Content Extensions
Boilerplate Extensions. We find another cluster comprised of extensions with no functionality released between 2018 and 2022. These extensions have identical dummy background pages and a browser action that loads the example HTML document from Listing 1. Looking at this document, we notice that the extensions are made using Extensionizr, a now-defunct web application for generating boilerplate browser extensions in a few clicks (Wolkov, 2023). While the extensions themselves do nothing, their store listings appear to promote blogs or businesses, with the featured image usually being a screenshot of the promoted website. As such, we posit these infringing extensions are just dummy items designed to take advantage of the CWS to advertise a given site.
@@ 6.3. Spam Extensions
One example is Cluster #16, which consists of extensions only declaring a browser action in their manifest. When clicking the extension icon, it loads an HTML document that has a link to the spammed website. These extensions rely on popular apps and services to grab the attention of CWS users, with names such as “123movie - 100% WORKING”, “Free Dogecoin Faucet” and “Whatsapp Plus APK”. Notably, “GS Auto Clicker:Free Download” is the only extension labeled as Malware in this cluster,333 https://crxcavator.io/report/chjhnkfpbgcajkfidohljkjjlfcmnahi which differs from the rest only in the link it promotes. Cluster #63 is also made up of extensions that have a browser action, but they take it a step further to give the appearance of having some functionality. This is achieved by having an HTML form with a few fields that relies on a simple JavaScript code to produce some—usually meaningless—output. This form also contains a call-to-action button promoting the spammed website. For example, the extension “arsenal vs everton live” has a football match score predictor that gives a random pair of numbers every time the form is submitted (mbscore.tv, 2023).
@@
Unreliable Clusters. We find 2 other clusters resembling those containing spam extensions, as they also have very similar manifests that just declare a browser action. While these clusters are mostly comprised of extensions with no functionality and links to external websites, they also contain low-effort extensions that perform an extremely simple task (e.g., a calculator). Unlike spam extensions, items in unreliable clusters lack a distinct set of API calls that can be used to group them together, and so they end up in different clusters than the former. We acknowledge that this is a limitation of our pipeline as currently designed, since extensions with short manifests and no API calls are clustered together. We discuss this point further in Section 8.
6.4. Other Extensions
@@
Chrome Apps. Apps are a legacy type of extension that were deprecated in 2020 and are currently only supported on ChromeOS devices (Google Chrome Developers, 2012b). They can be as simple as a shortcut that opens a URL in a new tab, being essentially an icon and a manifest that declares the app.launch.web_url property. As such, neither our tool (nor any tool) can determine whether they violate any policies without analysis of the linked external site.
Kiosk Apps. Cluster #19 is formed by kiosk apps generated using the official Chrome App Builder extension (Google, Inc., 2018). Akin to Chrome Apps, these extensions just load an external URL, but are intended to run full-screen on a dedicated kiosk or signage device.
Repackaged APKs. ChromeOS devices can run Android apps through the App Runtime for Chrome (ARC) (Google Chrome Developers, 2014). With this runtime, developers can bundle Android Application Package (APKs) into browser extensions and run them on a Chromebook. Because our pipeline cannot extract features from native APKs, we cannot cluster these extensions any further.
@@
6.5. Labeling Consistency
As described in Section 3, the CWS assigns a label to vetted extensions based on the reason for the takedown. After characterizing the top infringing clusters, we notice some discrepancies in the labels assigned by Google to these vetted extensions. This raises the question of whether the CWS vetting labels are indeed related to the behavior of an extension and if the process is consistent. To assess this, we identify which are the most common features among vetted extensions per label and compare the groups. We take all vetted extensions in our dataset, group them by vetting label, and count the occurrences of features in each group; that is, we count how many vetted extensions with a particular label have a given feature. We initially conducted this experiment including NTEs. However, doing so added too much noise to our results due to the large number of vetted NTEs and the seeming inconsistency of their vetting labels, as previously mentioned in Section 6.1. As such, we repeated the experiment and put extensions with “theme”, “wallpaper” or “new tab” in their name in a separate group.
Table 2. Top most occurring features in vetted extensions grouped by label. NTEs reported separately, highlighted rows show features in any top 5.
@@ 6.5. Labeling Consistency
Table 2 shows the top most occurring features found in vetted extensions per vetting label. Overall, there is a considerable overlap between the feature sets of the reported groups, with 5 features consistently appearing in the top 10 rankings regardless of the vetting label. Furthermore, these repeated features are related to the storage, runtime and tabs extension APIs, which are widely used in unvetted extensions as well. Still, we notice some interesting aspects that are worth discussing.
According to Google, one of the key reasons for transitioning to Manifest V3 is its “higher security and privacy guarantees” (Li, 2023). We see that malware authors are already transitioning to it, with this group having the highest ratio of extensions using MV3 among the studied groups (40%). One of the top API calls used by extensions in the Minor Policy Violation group is browser.browserAction.onClicked. This method is used to add a listener that will run when the extension icon is clicked, typically to open a new tab or a browser action popup. This validates the evidence we discuss in Section 6.3. Regarding NTEs, we see that more than half of them declare an URL that will open in a new tab when the extension is uninstalled, a much higher ratio than for the other groups. We also find that the topSites permission is used almost exclusively by NTEs, making it a good indicator of this group across vetted extensions.
@@
7. Characterizing CWS Malware
Out of all vetted extensions, those labeled as malware are of significant importance. In this section, we study which malware families are published in the CWS and use SimExt to cluster them by their behavior (RQ3). We note that we do not claim to study the entire browser extension malware ecosystem, as some families may be distributed outside of the CWS by other means, such as sideloading (Toulas, 2023; Zahravi and Girnus, 2023; Royer, 2023). Instead, we focus on malicious extensions that target the CWS.
7.1. Malware Detection
@@ 7.2. Malware Families
Table 3. Families of malware extensions detected by engines in VirusTotal.
@@
Table 3 shows the aforementioned groups of extensions that are known to VirusTotal per family. Almost half of the malicious extensions have no suggested threat classification label, closely followed by a generic “Trojan” group. Other labels such as “BrowExt”, “BroExtension” or “ChromeX” seem to only indicate that the malware is a browser extension, without adding any further information about the malware family or behavior. The only three extensions with more informative classifications are “SaveProtect VPN”,444 https://crxcavator.io/report/dodnpoijjkmcmlhlelmggejhfocfjgfc which is labeled as Adware; “Fea KeyLogger”,555 https://crxcavator.io/report/fgkghpghjcbfcflhoklkcincndlpobja which advertises itself as a keylogger; and “Free YouTube Subscribers Generator”,666 https://crxcavator.io/report/fdfchfidjajpidpjilnlboncflgnjdda which has a link to an external website instead of having any JavaScript code. From these findings, we conclude that vendor labels for malicious browser extensions are extremely poor, and in most cases non-existent.
@@ 8. Discussion
Detection of Infringing Extensions. The findings outlined in Section 5 support the assumption that the current CWS vetting process lacks scalable tooling for effectively finding or clustering similar extensions. Given the estimate that 86% of infringing extensions found by our tool are republished extensions (i.e., items that the store has seen and taken down before), we recommend using not only blocklists but also behavioral features to flag similar extensions, as this can contribute to improve the vetting process. We believe that similarity search tools such as SimExt can be a valuable complement to the human factor, helping to find infringing content faster and more accurately.
Repeat Offenders. We find evidence that the CWS does not ban publishers with multiple vetted extensions. Since 11% of the published infringing extensions found by our tool come from repeat offenders, we believe the CWS should enforce their own Repeat Abuse policy (Google Chrome Developers, 2022d) and suspend the accounts of repeat offenders.
Lifetime of Infringing Extensions. Infringing content stay for too long in the store, with 59% of vetted extensions remaining published for more than a year before they are taken down. Even for malware extensions, the median removal time is longer than 9 months at best. We consider these to be unacceptably high and believe that the CWS needs to focus its efforts on significantly reducing them. The addition of automated tools for finding items similar to previously vetted extensions should also help reduce these lifetimes.
Unpublished Infringing Extensions. As presented in Section 5, we find that some publishers of infringing extensions voluntarily remove them before receiving a takedown from Google. Given that repeat offenders are fairly common (and thus there is little repercussion for having multiple vetted extensions), we do not believe that these developers are acting like this to protect their accounts. Extensions labeled by Google as malware are remotely and automatically uninstalled from users browsers (Dunk, 2023). However, we find that extensions that are unpublished before being detected as malware are not subject to this process, leaving users compromised. Thus, we recommend that the CWS also assigns vetting labels to recently unpublished extensions as well as to vetted ones.
Vetting Labels. Providing labels for vetted content is a valuable transparency feature of the CWS vetting process. However, the current labeling scheme admits multiple improvements. One would be to move towards finer-grained labels that are more specific about the reasons for taking down the item, as these labels could better inform ecosystem studies such as ours. As mentioned in Section 6.1, more consistency when assigning labels to vetted extensions will also help increase confidence in the vetting process.
@@
8.1. Limitations and Future Work
@@
In Section 6.3, we discuss how our pipeline clusters together extensions with very few features, to the point that items with no logged API calls and different behaviors are difficult to distinguish from one another. Ideally, these unreliable clusters should be split further into as many sub-clusters as unique sets of behaviors. As an improvement, we propose expanding the API calls extracted by our static and dynamic analyses to include other frequently used objects, such as document and jQuery. This way, extensions with few or no calls to Extension APIs or inside the navigator object will be better clustered due to the use of more robust embeddings.
9. Related Work
@@
Previous research has crawled the CWS for various purposes related to the analysis of browser extensions (Kapravelos et al., 2014; Weissbacher et al., 2017; Aggarwal et al., 2018; Chen and Kapravelos, 2018; Pantelaios et al., 2020; Fass et al., 2021; Laperdrix et al., 2021; Solomos et al., 2022; Eriksson et al., 2022; Picazo-Sanchez et al., 2022).
@@
Extension Analysis. Pantelaios et al. statically extracted API calls to cluster extension version deltas to detect malicious updates (Pantelaios et al., 2020). Fass et al. used static analysis to find suspicious data flows in vulnerable extensions (Fass et al., 2021). Kapravelos et al. proposed a dynamic analysis architecture for identifying malicious behavior in extensions using on-the-fly generated pages (Kapravelos et al., 2014). Jagpal et al. used both static and dynamic analysis to capture behavioral signals from browser extensions (Jagpal et al., 2015), followed by Aggarwal et al. a few years later (Aggarwal et al., 2018). Picazo-Sanchez et al. used static, manual and dynamic analysis to mark malicious extensions (Picazo-Sanchez et al., 2022). Eriksson et al. combined both static and dynamic analysis to discover code vulnerabilities, uncovering the existence of NTEs that stole browsing traffic (Eriksson et al., 2022). Pantelaios and Kapravelos used dynamic execution to validate automated conversions of extensions from MV2 to MV3 (Pantelaios and Kapravelos, 2024). Unlike our work, previous dynamic analysis efforts have relied on running the extensions on an actual web browser through manual interaction or using tools like Selenium, Puppeteer and Playwright. Instead, we created a mocked V8 sandbox for faster code execution at scale.
@@
Finding Malicious Extensions with Machine Learning. Jagpal et al. used machine learning to flag malicious extensions based on behavioral signals, having to train a proprietary model daily to account for newly vetted extensions (Jagpal et al., 2015). Aggarwal et al. improved upon this by feeding sequences of API calls to their own Recurrent Neural Network (RNN) (Aggarwal et al., 2018). Pantelaios et al. used DBSCAN to cluster custom-made API sequences referred to as seeds (Pantelaios et al., 2020). Similarly, Picazo-Sanchez et al. used time series analysis and machine learning to cluster malicious extensions based on their download patterns (Picazo-Sanchez et al., 2022). In contrast to previous work, our approach does not try to classify extensions as benign or malicious, but to cluster similar extensions together. By using NLP and ZSL, we avoid the limitation of having to retrain our model when a new malicious behavior is discovered.
@@ Appendix A Serialization Example
Figure 7 provides an example of the feature serialization pipeline we describe in Section 4.2. The process takes as input the manifest key-value pairs of the extension and the API calls extracted by the static and dynamic analyzers. These items are converted to sentences that can then be merged into a single document.
Figure 7. Feature serialization example.

### v2

@@ VI-B No Content Extensions
Boilerplate Extensions. We find another cluster comprised of extensions with no functionality released between 2018 and 2022. These extensions have identical dummy background pages and a browser action that loads the example HTML document from Listing 1. Looking at this document, we notice that the extensions are made using Extensionizr, a now-defunct web application for generating boilerplate browser extensions in a few clicks [67]. While the extensions themselves do nothing, their store listings appear to promote blogs or businesses, with the featured image usually being a screenshot of the promoted website. As such, we posit these infringing extensions are just dummy items designed to take advantage of the CWS to advertise a given site.
@@ VI-C Spam Extensions
One example is cluster #16, which consists of extensions only declaring a browser action in their manifest. When clicking the extension icon, it loads an HTML document that has a link to the spammed website. These extensions rely on popular apps and services to grab the attention of CWS users, with names such as “123movie - 100% WORKING”, “Free Dogecoin Faucet” and “Whatsapp Plus APK”. Notably, “GS Auto Clicker:Free Download” is the only extension labeled as Malware in this cluster,444 https://crxcavator.io/report/chjhnkfpbgcajkfidohljkjjlfcmnahi which differs from the rest only in the link it promotes. Cluster #63 is also made up of extensions that have a browser action, but they take it a step further to give the appearance of having some functionality. This is achieved by having an HTML form with a few fields that relies on a simple JavaScript code to produce some—usually meaningless—output. This form also contains a call-to-action button promoting the spammed website. For example, the extension “arsenal vs everton live” has a football match score predictor that gives a random pair of numbers every time the form is submitted [35].
@@
Unreliable Clusters. We find 2 other clusters resembling those containing spam extensions, as they also have very similar manifests that just declare a browser action. While these clusters are mostly comprised of extensions with no functionality and links to external websites, they also contain low-effort extensions that perform an extremely simple task (e.g., a calculator). Unlike spam extensions, items in unreliable clusters lack a distinct set of API calls that can be used to group them together, and so they end up in different clusters than the former. We acknowledge that this is a limitation of our pipeline as currently designed, since extensions with short manifests and no API calls are clustered together. We discuss this point further in Section VIII.
VI-D Other Extensions
@@
Chrome Apps. Apps are a legacy type of extension that were deprecated in 2020 and are currently only supported on ChromeOS devices [11]. They can be as simple as a shortcut that opens a URL in a new tab, being essentially an icon and a manifest that declares the app.launch.web_url property. As such, neither our tool (nor any tool) can determine whether they violate any policies without analysis of the linked external site.
Kiosk Apps. Cluster #19 is formed by kiosk apps generated using the official Chrome App Builder extension [24]. Akin to Chrome Apps, these extensions just load an external URL, but are intended to run full-screen on a dedicated kiosk or signage device.
Repackaged APKs. ChromeOS devices can run Android apps through the App Runtime for Chrome (ARC) [12]. With this runtime, developers can bundle Android Application Package (APKs) into browser extensions and run them on a Chromebook. Because our pipeline cannot extract features from native APKs, we cannot cluster these extensions any further.
@@
VI-E Labeling Consistency
As described in Section III, the CWS assigns a label to vetted extensions based on the reason for the takedown. After characterizing the top infringing clusters, we notice some discrepancies in the labels assigned by Google to these vetted extensions. This raises the question of whether the CWS vetting labels are indeed related to the behavior of an extension and if the process is consistent. To assess this, we identify which are the most common features among vetted extensions per label and compare the groups. We take all vetted extensions in our dataset, group them by vetting label, and count the occurrences of features in each group; that is, we count how many vetted extensions with a particular label have a given feature. We initially conducted this experiment including NTEs. However, doing so added too much noise to our results due to the large number of vetted NTEs and the seeming inconsistency of their vetting labels, as previously mentioned in Section VI-A. As such, we repeated the experiment and put extensions with “theme”, “wallpaper” or “new tab” in their name in a separate group.
TABLE II: Top most occurring features in vetted extensions grouped by label. NTEs reported separately, highlighted rows show features in any top 5.
@@ VI-E Labeling Consistency
Table II shows the top most occurring features found in vetted extensions per vetting label. Overall, there is a considerable overlap between the feature sets of the reported groups, with 5 features consistently appearing in the top 10 rankings regardless of the vetting label. Furthermore, these repeated features are related to the storage, runtime and tabs extension APIs, which are widely used in published (non-vetted) extensions as well. Still, we notice some interesting aspects that are worth discussing.
According to Google, one of the key reasons for transitioning to Manifest V3 is its “higher security and privacy guarantees” [33]. We see that malware authors are already transitioning to it, with this group having the highest ratio of extensions using MV3 among the studied groups (40%). One of the top API calls used by extensions in the Minor Policy Violation group is browser.browserAction.onClicked. This method is used to add a listener that will run when the extension icon is clicked, typically to open a new tab or a browser action popup. This validates the evidence we discuss in Section VI-C. Regarding NTEs, we see that more than half of them declare a URL that will open in a new tab when the extension is uninstalled, a much higher ratio than for the other groups. We also find that the topSites permission is used almost exclusively by NTEs, making it a good indicator of this group across vetted extensions.
@@
VII Characterizing CWS Malware
Out of all vetted extensions, those labeled as malware are of significant importance. In this section, we study which malware families are published in the CWS and use SimExt to cluster them by their behavior (RQ3). We conclude with an analysis of malicious extensions with a high number of installs that are not present in VirusTotal. We note that we do not claim to study the entire browser extension malware ecosystem, as some families may be distributed outside of the CWS by other means, such as sideloading [62, 68, 55].
VII-A Malware Detection
@@ VII-B Malware Families
TABLE III: Families of malware extensions detected by engines in VirusTotal.
@@
Table III shows the aforementioned groups of extensions that are known to VirusTotal per family. Almost half of the malicious extensions have no suggested threat classification label, closely followed by a generic “Trojan” group. Other labels such as “BrowExt”, “BroExtension” or “ChromeX” seem to only indicate that the malware is a browser extension, without adding any further information about the malware family or behavior. The only three extensions with more informative classifications are “SaveProtect VPN”,555 https://crxcavator.io/report/dodnpoijjkmcmlhlelmggejhfocfjgfc which is labeled as Adware; “Fea KeyLogger”,666 https://crxcavator.io/report/fgkghpghjcbfcflhoklkcincndlpobja which advertises itself as a keylogger; and “Free YouTube Subscribers Generator”,777 https://crxcavator.io/report/fdfchfidjajpidpjilnlboncflgnjdda which has a link to an external website instead of having any JavaScript code. From these findings, we conclude that vendor labels for malicious browser extensions are extremely poor, and in most cases non-existent.
@@
VII-C Case Studies
Since almost all malware extensions flagged by the CWS are unknown to detection engines, we manually inspect those not present in VirusTotal and with at least 100k installs at the time of removal. Of the resulting 90 items that meet these criteria, we could not find a clear motivation for labeling 29 of them as malware.
A large group of the remaining extensions contain extensions with embedded tracking capabilities or that load remotely hosted code. One interesting group is formed by 12 extensions that use Google Tag Manager to download and execute an obfuscated script888 https://storage.googleapis.com/g1analytics/cloud_new_noab-obf.js from a Google Cloud Storage bucket. When unpacked, this script sends an HTTP request to a remote endpoint to get the country of the user and starts exfiltrating all visited URLs in real time if the location matches the United States. To avoid detection, the payload of these requests is encoded as a binary blob. Some extensions that request the webRequestBlocking permission load an additional remote script,999 https://storage.googleapis.com/analytics-cloud/js_analytics_protected.js which intercepts all search queries from a predefined list of lesser known search engines and redirects them to another domain.
We also find 4 trojanized extensions that run code in Facebook to obtain authentication tokens (such as fb_dtsg). The intercepted tokens are then used to make requests to the GraphQL endpoint in the background without the user’s awareness. Once authenticated, the user is joined to a Facebook group that varies depending on the extension, and the device’s access token is posted there. Along with the token, these extensions also exfiltrate the number of business accounts associated with that credential, presumably to triage victims.
Appendix C lists the extensions discussed above.
@@ VIII Discussion
Detection of Infringing Extensions. The findings outlined in Section V support the assumption that the current CWS vetting process lacks scalable tooling for effectively finding or clustering similar extensions. Given the estimate that 86% of infringing extensions found by our tool are republished extensions (i.e., items that the store has seen and taken down before), we recommend using not only blocklists but also behavioral features to flag similar extensions, as this can contribute to improve the vetting process. We believe that similarity search tools such as SimExt can be a valuable complement to the human factor, helping to find infringing content faster and more accurately.
Repeat Offenders. We find evidence that the CWS does not ban publishers with multiple vetted extensions. Since 11% of the published infringing extensions found by our tool come from repeat offenders, we believe the CWS should enforce their own Repeat Abuse policy [19] and suspend the accounts of repeat offenders.
Lifetime of Infringing Extensions. Infringing content stays for too long in the store, with 59% of vetted extensions remaining published for more than a year before they are taken down. Even for malware extensions, the median removal time is longer than 9 months at best. We consider these to be unacceptably high and believe that the CWS needs to focus its efforts on significantly reducing them. The addition of automated tools for finding items similar to previously vetted extensions should also help reduce these lifetimes.
Unpublished Infringing Extensions. As presented in Section V, we find that some publishers of infringing extensions voluntarily remove them before receiving a takedown from Google. Given that repeat offenders are fairly common (and thus there is little repercussion for having multiple vetted extensions), we do not believe that these developers are acting like this to protect their accounts. Extensions labeled by Google as malware are remotely and automatically uninstalled from users browsers [7]. However, we find that extensions that are unpublished before being detected as malware are not subject to this process, leaving users compromised. Thus, we recommend that the CWS also assigns vetting labels to recently unpublished extensions as well as to vetted ones.
Vetting Labels. Providing labels for vetted content is a valuable transparency feature of the CWS vetting process. However, the current labeling scheme admits multiple improvements. One would be to move towards finer-grained labels that are more specific about the reasons for taking down the item, as these labels could better inform ecosystem studies such as ours. As mentioned in Section VI-A, more consistency when assigning labels to vetted extensions will also help increase confidence in the vetting process.
@@
VIII-A Limitations and Future Work
@@
In Section VI-C, we discuss how our pipeline clusters together extensions with very few features, to the point that items with no logged API calls and different behaviors are difficult to distinguish from one another. Ideally, these unreliable clusters should be split further into as many sub-clusters as unique sets of behaviors. As an improvement, we propose expanding the API calls extracted by our static and dynamic analyses to include other frequently used objects, such as document and jQuery. This way, extensions with few or no calls to Extension APIs or inside the navigator object will be better clustered due to the use of more robust embeddings.
IX Related Work
@@
Previous research has crawled the CWS for various purposes related to the analysis of browser extensions [29, 66, 1, 3, 48, 9, 31, 58, 8, 51, 44].
@@
Extension Analysis. Pantelaios et al. statically extracted API calls to cluster extension version deltas to detect malicious updates [48]. Fass et al. used static analysis to find suspicious data flows in vulnerable extensions [9]. Kapravelos et al. proposed a dynamic analysis architecture for identifying malicious behavior in extensions using on-the-fly generated pages [29]. Jagpal et al. used both static and dynamic analysis to capture behavioral signals from browser extensions [28], followed by Aggarwal et al. a few years later [1]. Picazo-Sanchez et al. used static, manual and dynamic analysis to mark malicious extensions [51]. Eriksson et al. combined both static and dynamic analysis to discover code vulnerabilities, uncovering the existence of NTEs that stole browsing traffic [8]. Pantelaios and Kapravelos used dynamic execution to validate automated conversions of extensions from MV2 to MV3 [47]. Unlike our work, previous dynamic analysis efforts have relied on running the extensions on an actual web browser through manual interaction or using tools like Selenium, Puppeteer and Playwright. Instead, we created a mocked V8 sandbox for faster code execution at scale.
Defining Potentially Harmful Extensions. Somé was the first to propose a taxonomy of sensitive APIs to define extensions that pose a security or privacy risk to the user [59]. Hsu et al. introduced the concept of Security-Noteworthy Extensions to extend this set to cover malware, vulnerable, and policy-violating extensions [26]. In our work, we focus on evaluating the CWS vetting process by using its own vetting labels, rather than seeking to define what potentially harmful extensions are.
@@
Finding Malicious Extensions with Machine Learning. Jagpal et al. used machine learning to flag malicious extensions based on behavioral signals, having to train a proprietary model daily to account for newly vetted extensions [28]. Aggarwal et al. improved upon this by feeding sequences of API calls to their own Recurrent Neural Network (RNN) [1]. Pantelaios et al. used DBSCAN to cluster custom-made API sequences referred to as seeds [48]. Similarly, Picazo-Sanchez et al. used time series analysis and machine learning to cluster malicious extensions based on their download patterns [51]. In contrast to previous work, our approach does not try to classify extensions as benign or malicious, but to cluster similar extensions together. By using NLP and ZSL, we avoid the limitation of having to retrain our model when a new malicious behavior is discovered.
@@ Appendix A Serialization Example
Figure 7 provides an example of the feature serialization pipeline we describe in Section IV-B. The process takes as input the manifest key-value pairs of the extension and the API calls extracted by the static and dynamic analyzers. These items are converted to sentences that can then be merged into a single document.
Figure 7: Feature serialization example.